OpenStudy (bshakoor):

What is the mean numbers of 320,280,280,250,270?

OpenStudy (evoker):

The mean is the sum which is then divided by the number of numbers.

OpenStudy (astrophysics):

In other words the average

OpenStudy (mathmale):

Add together the numbers in this list: 320,280,280,250,270. How many numbers are there in the list? Divide the sum by the number of numbers in the list (which is 5). Show your work.

OpenStudy (mathmale):

Please share your work. If you're already done with this problem, would you please close this post by clicking on the "Close" button, above. Thank you.

OpenStudy (ppery329):

When you add 320, 280, 280, 250 and 270, you would get 1400. 280+280=560+320=880+250=1130+270=1400. 1400/5=280. The mean of 320, 280, 280, 250 and 270 is 280.